A ball with mass of 0.050 kg is dropped from a height of 2.0 m. It collides with the floor, then sticks to the table. The Collison takes 0.020 s. a) Use the conservation of mechanical energy (Ei = Ef) to calculate the velocity of the ball before collision with the floor
Answer #1

a)
intially ball have only potential energy = m*g*h
finally it has only kinetic energy = 0.5*m*v^2

so, use
Ei = Ef
m*g*h = 0.5*m*v^2
g*h = 0.5*v^2
9.8*2.0 = 0.5*v^2
v^2 = 39.2
v = 6.26 m/s

Answer: 6.26 m/s